OpenStudy (anonymous):

Ana is a girl and has the same number of brothers as sisters. Andrew is a boy and has twice as many sisters as brothers. Ana and Andrew are the children of Emma. How many children does Emma have? (A) 2 (B) 3 (C) 5 (D) 7 (E) 8

OpenStudy (anonymous):

@hope_nicole

OpenStudy (anonymous):

follow @Luis_Rivera steps, they are very good

OpenStudy (anonymous):

emma has 4 female children, and the male children are n − 1 = 4 − 1 = 3. total number of children Emma has is 4 + 3 = 7. The answer is D

OpenStudy (anonymous):

cool, thkns